I have an exam coming up and i've completely blanked out. Could you give me an example on how to solve 6x + 9y = 15 ? Thank you so much!
Answer by xcentaur(357)

Since you have 2 variables x and y in the given equation, you would require one more equation to obtain values of x and y.


6x + 9y = 15
3[2x + 3y = 5]


2x + 3y = 5


then,
x=(5-3y)/2
y=(5-2x)/3


Substitute either of these two relative values of x or y in the 2nd equation.
Then you will have an equation entirely in x or y.
solve to obtain the required value,and then use the relative value to obtain the second variable.


Hope this helped,
Best of luck.